Problem 7-18

Assume Gillette Corporation will pay an annual dividend of $0.65 one year from now. Analysts expect this dividend to grow at 12.0% per year thereafter until the 5th year. Thereafter, growth will level off at 2.0% per year. According to the​ dividend-discount model, what is the value of a share of Gillette stock if the​ firm's equity cost of capital is 8.0%​?
Answer #1

Dividend in year 1=0.65
Dividend in year 2=0.65*(1+Growth rate)=0.65*(1+12.0%)=0.728
Dividend in year 3=0.65*(1+12.0%)^2=0.81536
Dividend in year 4=0.65*(1+12.0%)^3=0.9132032
Dividend in year 5=0.65*(1+12.0%)^4=1.022787584

Terminal value=(Dividend in year 5)*(1+Constant growth rate)/(Cost of capital-Constant growth rate)
Terminal value=(1.022787584)*(1+2.0%)/(8.0%-2.0%)
Terminal value=(1.043243336)/(6%)=17.38738893


Value of share=Present value of future cash flows
=0.65/(1+8%)^1+0.728/(1+8%)^2+0.81536/(1+8%)^3+0.9132032/(1+8%)^4+1.022787584/(1+8%)^5+17.38738893/(1+8%)^5
=0.601851852+0.624142661+0.647259056+0.671231614+0.696092044+11.83356475
=15.07414 or $15.07 (Rounded to 2 decimal places)